OpenStudy (anonymous):

4 times a number plus 4 is twice the number

OpenStudy (anonymous):

4x+4 = 2x 2x = -4 x = -2 which is the number

OpenStudy (lena772):

4x+4=2x

OpenStudy (lena772):

ok juststoper beat me to it

OpenStudy (anonymous):

thannkkk youuu (:

OpenStudy (doc.brown):

\[4x+4=2x\]\[4x-2x+4=2x-2x\]\[2x+4=0\]\[2x+4-4=0-4\]\[2x=-4\]\[\frac{2x}{2}=\frac{-4}{2}\]\[1x=-2\]\[x=-2\]